OpenStudy (anonymous):

Given that f(x) = x2 − 3x + 3 and g(x) = the quantity of x minus one, over four, solve for f(g(x)) when x = 5. A) -3 B) 1 C)7 D)13

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

\[\large f(x) = x^2 - 3x + 3\] \[\large g(x) = \frac{x - 1}{4}\] So what we do...is take our inside function *g(x)* and solve it when x = 5 can you do that?

OpenStudy (anonymous):

okay so 5-1/4

OpenStudy (johnweldon1993):

mmhmm :) which is?

OpenStudy (anonymous):

4/4 which is obvisously 1

OpenStudy (johnweldon1993):

Indeed...so now..we take that result (1) and we plug that into f(x) wherever we see an 'x' \[\large f(1) = 1^2 - 3(1) + 3\]